Two dozen units in an apartment complex need to be painted. It takes 3 gallons of paint to cover each apartment.
How many apartments can be painted with 36 gallons?
How many gallons will it take to paint all 24 apartments?

Answer by gonzo(654)

24 units need to be painted.
3 gallons of paint each.
-----
you would need 24*3 gallons of paint to cover them all.
24 * 3 = 72 gallons required to paint all 24.
-----
you can paint 36/3 apartments with 36 gallons.
36 / 3 = 12 apartments that can be painted with 36 gallons.
-----
your general formula would be:
let a = number of apartments
let g = number of gallons required to paint each apartment.
let t = total number of gallons available.
-----
a * g = t
meaning the number of apartments times the number of gallons required for each equals the total number of gallons required.
-----
t / g = a
meaning the total number of gallons available divided by the number of gallons required for each apartment equals the number of apartments that can be painted.
-----